A firm wants to minimize cost when Q=160=2√KL ;r=4 k=2
find K and L




1
Expert's answer
2021-10-01T16:54:22-0400

"2 \sqrt{KL}=160 \\

L = 2l+4K-λ[\sqrt{KL} -80] \\

\frac{∂L}{∂l} = 2 - \frac{λ}{2} \sqrt{\frac{K}{l}} =0 \\

λ = 4 \sqrt{\frac{l}{K}} \\

\frac{∂L}{∂K} = 4 -\frac{λ}{2} \sqrt{\frac{K}{l}} \\

l = 2K \\

2 \sqrt{2} K =160 \\

K=56.58 \\

L=113.15"